OpenStudy (anonymous):

need help solve this equation using the elimination method 6x – 4y = 54 –9x + 2y = –69

OpenStudy (ajprincess):

Multiply the second equation by 2 nd can u plz tell me vat u get?

OpenStudy (anonymous):

-18+4y=-138

OpenStudy (ajprincess):

yup good:) nw add this one with the first equation. terms with variable y will be elimnated. then solve for x. can u do it @yayasims21

OpenStudy (anonymous):

i cannt can u do it

OpenStudy (ajprincess):

plz give it a try. if u face any difficulty when doing it plz ask. I will help u:)

OpenStudy (anonymous):

ok so im suppose to add -18+4y=-138 to 6x-4y=54

OpenStudy (ajprincess):

ya:)

OpenStudy (anonymous):

i cant do it my calculator wont let me

OpenStudy (ajprincess):

-18x+4y+6x-4y=-138+54 (-18x+6x)+(4y-4y)=-138+54 See if u can do it nw:)

OpenStudy (ajprincess):

Add the ones inside the parentheses

OpenStudy (anonymous):

i cant do it i really suck at math

OpenStudy (ajprincess):

what is -18x+6x?

OpenStudy (anonymous):

-12

OpenStudy (ajprincess):

-12x. what is 4y-4y?

OpenStudy (anonymous):

0

OpenStudy (ajprincess):

yup right:) what is -138+54

OpenStudy (anonymous):

-84

OpenStudy (ajprincess):

yup good. nw combine these results.

OpenStudy (anonymous):

add them

OpenStudy (ajprincess):

ya.

OpenStudy (anonymous):

i get -96

OpenStudy (theviper):

@ajprincess happy new year & to @yayasims21 too :)

OpenStudy (anonymous):

o thanks happy new year to u 2

OpenStudy (theviper):

:)

OpenStudy (ajprincess):

no. remember u can add -12x with -84. nd -84 is in the right side of the equal aign. -18x+6x+4y-4y=-138+54 -12x+0=-84 -12x=-84 Divide by -12 both sides to find x. Thanx @TheViper and I wish u the same:)

OpenStudy (theviper):

:)

OpenStudy (anonymous):

-12 divided by -12 is 1 and -84 is 7

OpenStudy (ajprincess):

remember u r dividing -12x by -12 nt -12. this leaves u with x and u r right abt getting 7:)

OpenStudy (ajprincess):

So ur x becomes 7. To find y plug in this value in either 6x – 4y = 54 or –9x + 2y = –69 and solve

OpenStudy (anonymous):

ok wat do u mean how am i suppose to do tht

OpenStudy (ajprincess):

x=7. Plug this in 6x-4y=54 6(7)-4y=54 42-4y=54 1.Subtract 42 both sides. 2.Divide by -4 both sides to find y.

OpenStudy (anonymous):

i cant do this im so confused

OpenStudy (ajprincess):

42-4y=54 1.Subtract 42 both sides. 2.Divide by -4 both sides to find y.

OpenStudy (ajprincess):

Plz follow the steps

OpenStudy (anonymous):

i am but so im suppose to subtract 42from -9

OpenStudy (ajprincess):

no. let me do one step and show. 42-4y=54 Subtract 42 both sides 42-42-4y=54-42 0-4y=12 -4y=12 Nw divide by -4 both sides to find y

OpenStudy (anonymous):

-3x

OpenStudy (ajprincess):

-4y=12 Divide by -4 both sides (-4y)/-4=12/-4 y=-3

OpenStudy (anonymous):

ok i got it is there anymore we have to do

OpenStudy (ajprincess):

nope. evrythng is done:)

OpenStudy (anonymous):

ok well can u show this all at once so i can write it down

OpenStudy (anonymous):

can you @ajprincess

OpenStudy (ajprincess):

Extremely sorrry @yayasims21. busy right nw.

OpenStudy (anonymous):

o alright illtry to put it together

OpenStudy (ajprincess):

hav u finished @yayasims21?

OpenStudy (ajprincess):

\(6x-4y=54-(1)\) \(-9x+2y=-69-(2)\) (2)*2=>\(-18x+4y=-138-(3)\) (3)+(1)=>\(-18x+4y+6x-4y=-138+54\) \(-12x=-84\) \(-12x/-12=-84/-12\) \(x=7\) Substituting this in (1)=> \(6(7)-4y=54\) \(42-4y=54\) \(-4y=54-42\) \(-4y=12\) \(-4y/-4=12/-4\) \(y=-3\)